OpenStudy (anonymous):

Solve system of equations. 2x-4y=10 7x-9y=12 Don't understand this at all.

OpenStudy (merchandize):

with the help of one of these equation take out the value of x or y. for example 2x=10+ 4y. and then x=10+4y/2. now put the value of x in other equation and then find out the value of x and y by solving the equation.

OpenStudy (texaschic101):

2x + 4y = 10 2x = 10 - 4y x = 10/2 - 4/2y x = 5 - 2y now sub 5 - 2y in for x in the other equation. Do you understand what I did so far ?

OpenStudy (anonymous):

Yes thank you

OpenStudy (texaschic101):

now you can solve for y by subbing in 5 - 2y in for x. Can you do that ?

OpenStudy (anonymous):

Yes I got 3/5,-11/5

OpenStudy (texaschic101):

let me check.... 7x - 9y = 12 7(5 - 2y) - 9y = 12 (distribute through the parenthesis) 35 - 14y - 9y = 12 (combine like terms) 35 - 23y = 12 35 - 12 = 23y 23 = 23y y = 1 now sub 1 in for y 2x - 4y = 10 2x - 4(1) = 10 2x -4 = 10 2x = 10 + 4 2x = 14 x = 7 check... 2x - 4y = 10 2(7) - 4(1) = 10 14 - 4 = 10 10 = 10 (correct) Something might be wrong because when I check it with the other equation, it doesn't equal......

OpenStudy (texaschic101):

@merchandize .....what did I do wrong...it checks out with one equation, but I can't get it to equal with the other equation when I check it.

OpenStudy (texaschic101):

@burhan101 ...can you come take a look at this please

OpenStudy (anonymous):

I have 3 answer choices (3/5,-11/5) (-21/5, -23/5) (-69/5,-47/5)

OpenStudy (texaschic101):

Let me do it the elimination way...hold on

OpenStudy (texaschic101):

2x - 4y = 10 -->(7)2x - 4y = 10 7x - 9y = 12 --> (-2)7x - 9y = 12 --------------- 14x - 28y = 70 (result of multiplying by 7) -14x + 18y = - 24 (result of multiplying by -2) ------------ - 10y = 46 y = - 46/10 = - 23/5 the answer is going to be (-21/5, -23/5)

OpenStudy (merchandize):

(-21/5,-23/5) is the answer.